# Likelihood Based Inference for ARMA Models

### v5

@@ Abstract
Autoregressive moving average (ARMA) models are widely used for analyzing time series data. However, standard likelihood-based inference methodology for ARMA models has avoidable limitations. We show that common ARMA likelihood maximization strategies often lead to sub-optimal parameter estimates. While this possibility has been previously identified, no routinely applicable algorithm has been developed to resolve the issue. We introduce a novel random initialization algorithm, designed to take advantage of the structure of the ARMA likelihood function, which overcomes these optimization problems. Additionally, we show that profile confidence intervals provide superior confidence intervals to those based on the Fisher information matrix. The efficacy of the proposed methodology is demonstrated through a data analysis example and a series of simulation studies. This work makes a significant contribution to statistical practice by identifying and resolving under-recognized shortcomings of existing procedures that frequently arise in scientific and industrial applications.
Keywords: Auto-regressive, Moving average, Maximum likelihood, Profile likelihood, Optimization
[…]
algorithm
@@ 1 Introduction
Auto-regressive moving average (ARMA) models are the most well known and frequently used approach to modeling time series data. The general ARMA model was first described by Whittle (1951), and popularized by Box and Jenkins (1970). Today, ARMA models are a fundamental component of various academic curricula, leading to their widespread use in both academia and industry. ARMA models are as foundational to time series analysis as linear models are to regression analysis, and they are often used in conjunction for regression with ARMA errors. A Google Scholar search for articles from 2024 onward that include the phrase “time series” (and variants), and the term “ARMA” (and variants) yields over 14,000 results. While not all these articles focus on the same models discussed here, the importance of this model class to modern science cannot be overstated. Given the ubiquity of ARMA models, even small improvements in parameter estimation constitute a significant advancement of statistical practice.
@@
Imperfect likelihood optimization has an immediate consequence of complicating standard model selection procedures. Algorithms in widespread use lead to frequent inconsistencies in which a smaller model is found to have a higher maximized likelihood than a larger model within which it is nested. This is mathematically impossible but can occur when the likelihood is imperfectly maximized, and is commonly observed using contemporary methods for ARMA models. Such inconsistencies are a distraction for the interpretation of results even when they do not substantially change the conclusion of the data analysis. Removing numeric inconsistencies can, and should, increase confidence in the correctness of statistical inferences.
There are various software implementations available for the estimation of ARMA model parameters. In this article, we focus on the standard implementations in R (stats package) and Python (statsmodels module), which we selected due to their widespread usage. While both implementations offer multiple ways to estimate parameters, the default approach in both software packages is to perform likelihood maximization, assuming that the error terms are Gaussian. The challenges in parameter estimation arise in this situation because there is no closed-form expression for the likelihood function, though computational algorithms do exist for maximizing ARMA likelihoods (Gardner et al., 1980).
We begin by providing essential background information on the estimation of ARMA model parameters. We then present our proposed approach for parameter estimation, which leads to parameter values with a likelihood that is never lower and sometimes higher than the standard method. This is followed by a motivating example and a discussion of the potential implications of our proposed method. We also discuss the construction of standard errors for our maximum likelihood estimate. Specifically, we show that estimates of the standard error for model parameters that are default output of R and Python can be misleading, and we provide a reliable alternative. Throughout the article, we use the stats::arima function from the R programming language as a baseline for comparison. The same methodology for fitting parameters is used in the statsmodels.tsa module in Python, and we demonstrate in Appendix B that our approach applies to this software as well.
2 Maximum Likelihood for ARMA Models
Following the notation of Shumway and Stoffer (2017), a time series $\{x_{t};t=0,\pm 1,\pm 2,\ldots\}$ is said to be $\text{ARMA}(p,q)$ if it is (weakly) stationary and
@@ 2 Maximum Likelihood for ARMA Models
with $\{w_{t};t=0,\pm 1,\pm 2,\ldots\}$ denoting a mean zero white noise (WN) processes with variance $\sigma_{w}^{2}>0$, and $\phi_{p}\neq 0$, $\theta_{q}\neq 0$. We refer to the positive integers $p$ and $q$ of Eq. 1 as the autoregressive (AR) and moving average (MA) orders, respectively. A non-zero intercept could also be added to Eq. 1, but for simplicity we assume that the time series $\{x_{t};t=0,\pm 1,\pm 2,\ldots\}$ has zero mean. We denote the set of all model parameters as $\psi=\{\phi_{1},\ldots,\phi_{p},\theta_{1},\ldots,\theta_{q},\sigma_{w}^{2}\}$. Our objective is to estimate model parameters using the observed data.
Given the importance of ARMA models, numerous methods have been developed for parameter estimation. For instance, parameters can be estimated through methods such as Bayesian inference (Monahan, 1983; Chib and Greenberg, 1994) or neural networks (Chon and Cohen, 1997), among many others. Specialized methods include those for integer-valued data (Weiß and Zhu, 2024), approaches robust to outliers (Chakhchoukh, 2010), and non-Gaussian ARMA processes (Lii, 1990). While each methodology has its merits, we focus on the approach most widely adopted in statistical practice: likelihood maximization, under the assumption that the WN process $\{w_{t}\}$ is Gaussian.
@@
A natural estimator would involve minimizing the sum of squares $\sum_{t=1}^{n}w_{t}^{2}$. However, since only $x_{1},x_{2},\ldots,x_{n}$ are observed and $w_{t}$ is recursively defined in Eq. 2 using values of $x_{t-p}$ and $w_{t-q}$, directly minimizing this sum is intractable. The CSS method addresses this issue by conditioning on the first $p$ values of the process, assuming $w_{p}=w_{p-1}=\ldots=w_{p+1-q}=0$, and minimizing the conditioned sum $\sum_{t=p+1}^{n}w_{t}^{2}$. While the CSS method provides an attractive solution do to its relative simplicity and easiness to compute, it ignores the error terms for the first few observations. This is particularly concerning when the time series is short or when there are missing observations. CSS minimization was previously popular because methods for likelihood maximization were considered prohibitively slow, though this is no longer the case with currently available hardware and software (Ripley, 2002).
For likelihood maximization, Eq. 1 is reformulated as an equivalent state-space model. Although there are several ways this can be done, the approach of Gardner et al. (1980) is widely used. In this approach, we let $r=\max(p,q+1)$ and extend the set of parameters so that $\bar{\psi}=\{\phi_{1},\allowbreak\ldots,\allowbreak\phi_{r},\allowbreak\theta_{1},\allowbreak\ldots,\allowbreak\theta_{r-1},\allowbreak\sigma^{2}_{w}\}$, with some of the $\phi_{i}s$ or $\theta_{i}s$ being equal to zero unless $p=q+1$. We define a latent state vector $z_{t}\in\mathbb{R}^{r}$, along with transition matrices $T\in\mathbb{R}^{r\times r}$ and $Q\in\mathbb{R}^{r\times 1}$, enabling the recovery of the original sequence $\{x_{t}\}$ using Equations 3 and 4. For a detailed explanation of how to define the latent state $z_{t}$ and transition matrices $T$ and $Q$ in order to recover the ARMA model, we refer readers to Chapter 3 of Durbin and Koopman (2012). Along with initializations for the mean and variance of $z_{0}$, these equations allow for the exact computation of the likelihood of the ARMA model via the Kalman filter (Kalman, 1960), which can subsequently be optimized by a numeric procedure such as the BFGS algorithm (Fletcher, 2000).
@@
Numeric black-box optimizers require an initial guess for parameter values. For ARMA models, this task is non-trivial because the valid parameter region is defined by polynomial roots rather than the parameters themselves, as discussed in the next section. The default strategy in R and Python is to use the CSS estimator for initialization. This is an effective approach because the CSS estimator asymptotically converges to the MLE (Shumway and Stoffer, 2017), and may therefore be close to the global maximum when there are sufficiently many observations. However, the CSS initialization is less useful with limited data, or when there are missing observations. The CSS estimate may also lie outside the valid parameter region, and in such cases, parameters are reinitialized at the origin. Both software implementations also allow for manual selection of initial parameter values, but finding suitable initializations manually can be challenging due to complex parameter inter-dependencies.
@@
The log-likelihood function of ARMA models is often multimodal (Ripley, 2002), and therefore this single initialization approach can result in parameter estimates corresponding to local maxima (see Figure 1). This is true even for a carefully chosen initialization, such as the CSS estimate. A common strategy to optimize multimodal loss functions is to perform multiple optimizations using different initial parameters. However, we have found no instances of practitioners using a multiple initialization strategy for estimating ARMA model parameters. This may be explained by a general unawareness of the practical possibility of converging to a local maximum or because obtaining a suitable collection of initializations for ARMA models is nontrivial. For example, independently initializing parameters at random can place the parameter vector outside the region of interest, and can fail to adequately cover the plausible region (Appendix A).
@@
Figure 1: The profile log-likelihood of data simulated from four distinct MA(1) models, demonstrating a few examples of multimodal likelihood surfaces. The solid, black line indicates the true value of $\theta_{1}$; the dotted line is the CSS-initialization. The dashed lines correspond to the estimate $\hat{\theta}_{1}$ using stats:arima (red) and Algorithm 1 (implemented in arima2::arima, blue).
2.1 A Novel Multi-start Algorithm
@@
To obtain random parameter initializations, parameter sets must correspond to causal and invertible ARMA processes; definitions are in Chapter 3 of Shumway and Stoffer (2017). Let $\{\phi_{i}\}_{i=1}^{p}$ and $\{\theta_{i}\}_{i=1}^{q}$ be the coefficients of the $\text{ARMA}(p,q)$ model (Eq. 1), and define $\Phi(x)=1-\phi_{1}x-\phi_{2}x^{2}-\ldots-\phi_{p}x^{p}$ as the AR polynomial, and $\Theta(x)=1+\theta_{1}x+\theta_{2}x^{2}+\ldots+\theta_{q}x^{q}$ as the MA polynomial. An ARMA model is causal and invertible if the roots of the AR and MA polynomials lie outside the complex unit circle. Therefore in order to obtain valid, random parameter initializations, we sample the roots of $\Phi(x)$ and $\Theta(x)$ and use these values to reconstruct parameter initializations.
It is an easier task to sample inverted roots, as the sufficient conditions for causality and invertibility now require that the inverted roots lie inside the complex unit circle, a region easier to sample uniformly. The roots of the polynomials can be real or complex; complex roots must come in complex conjugate pairs in order for all of the corresponding model parameters to be real. The simplest approach would be to sample all inverted root pairs ($z_{1},z_{2}$) within the complex unit circle by uniformly sampling angles and radii (lines 10-14 of Algorithm 1). However, this would imply almost surely all root pairs are complex, and some model parameters would only be sampled as positive (or negative). For instance, consider an AR(2) model. The AR polynomial is:
In this equation, if both $z_{1},z_{2}$ are complex conjugates, then $\phi_{2}=z_{1}z_{2}>0$. As such, the only way that $\phi_{2}<0$ is if $z_{1},z_{2}\in\mathbb{R}$. Similar results hold for the MA coefficients, though the signs are flipped. This issue is directly addressed in lines 5-8 of Algorithm 1: root pairs are sampled as real with probability $p=\sqrt{1/2}$, and real pairs sampled with the same sign with probability $p$, such that the product (and sums) of each pair is positive with probability $1/2$. We sample conjugate pairs within an annular disk on the complex plane to avoid trivial and approximately non-stationary cases. The radii of both the inner and outer circles defining the disk are defined using $\gamma$ in lines 7, 10, and 13.
Parameter redundancy in ARMA models occurs when the polynomials $\Phi(x)$ and $\Theta(x)$ share one or more roots, leading to an overall reduction in model order. This complicates parameter initialization (Line 4 of Algorithm 1), optimization, and identifiability. The ARMA model (Eq. 1) can be rewritten as:
$\Phi(B)x_{t}=\Theta(B)w_{t},$ (5)
where $B$ is the backshift operator, i.e., $Bx_{t}=x_{t-1}$. Using the fundamental theorem of algebra, Equation 5 can be factored into $(1-\lambda_{1}B)\ldots(1-\lambda_{p}B)x_{t}=(1-\nu_{1}B)\ldots(1-\nu_{q}B)w_{t}$, where $\{\lambda_{i}\}_{i=1}^{p}$ and $\{\nu_{j}\}_{j=1}^{q}$ are the inverted roots of $\Phi(B)$ and $\Theta(B)$, respectively. If $\lambda_{i}=\nu_{j}$ for any $(i,j)\in\{1,\ldots,p\}\times\{1,\ldots,q\}$, then the roots will cancel each other out, resulting in an ARMA model of smaller order. As an elementary example, consider the $\text{ARMA}(1,1)$ and $\text{ARMA}(2,2)$ models in equations 6 and 7.
$x_{t}$ $=\frac{1}{3}x_{t-1}+w_{t}+\frac{2}{3}w_{t-1},$ (6)
$x_{t}$ $=\frac{5}{6}x_{t-1}-\frac{1}{6}x_{t-2}+w_{t}+\frac{1}{6}w_{t-1}-\frac{1}{3}w_{t-2},$ (7)
While these two models appear distinct at first glance, re-writing the models in polynomial form (Eq. 5) shows that these two models are actually equivalent after canceling out the common factors on each side of Eq. 5. In a similar fashion, it is possible that the roots are not exactly equal but are approximately equal. In this case, the ratio of factors becomes close to one, resulting in a similar effect to when the roots exactly cancel. We avoid the possibility of nearly canceling roots in parameter initializations by requiring the minimum Euclidean distance between inverted polynomial roots to be greater than $\alpha$. This is done in line 4 of Algorithm 1, though the condition is rarely triggered if the order of the model is of typical size $(p,q<4)$.
Our sampling scheme is combined with existing procedures for numeric optimization of model log-likelihoods (lines 1 and 16), as well as the default initialization strategy for $\psi_{0}$ used by existing software (such as the CSS initialization). Doing so guarantees that final estimates correspond to likelihood values greater than or equal to the currently accepted standards in the software environment where the algorithm is implemented. For this article, both the numeric optimization procedure $f(\cdot)$ and the parameter initialization strategy to obtain $\psi_{0}$ are those implemented in stats::arima. The stopping criterion was chosen so that the algorithm stops trying new initial values when no new maximum has been found using the last $M$ parameter initializations. Alternative stopping criterion can be used (see for example, Hart (1998)), but we found that this simple heuristic works well in practice.
Algorithm 1 is implemented in the R package arima2, available on the Comprehensive R Archive Network (CRAN) (Wheeler et al., 2023). The package features the function arima2::arima, which is an adaptation of the stats::arima function modified to incorporate the adjustments specified by Algorithm 1.
8 Get preliminary estimate: $\hat{\psi}_{0}=f(\psi_{0})$; set $k=0$;
@@
12           Sample paired roots as real with probability $p$;

### v6

# Revisiting Inference for ARMA Models: Improved Fits and Superior Confidence Intervals

@@ Abstract
Autoregressive moving average (ARMA) models are widely used for analyzing time series data. However, standard likelihood-based inference methodology for ARMA models has avoidable limitations. We show that currently accepted standards for ARMA likelihood maximization frequently lead to sub-optimal parameter estimates. Existing algorithms have theoretical support, but can result in parameter estimates that correspond to a local optimum. While this possibility has been previously identified, it remains unknown to most users, and no routinely applicable algorithm has been developed to resolve the issue. We introduce a novel random initialization algorithm, designed to take advantage of the structure of the ARMA likelihood function, which overcomes these optimization problems. Additionally, we show that profile likelihoods provide superior confidence intervals to those based on the Fisher information matrix. The efficacy of the proposed methodology is demonstrated through a data analysis example and a series of simulation studies. This work makes a significant contribution to statistical practice by identifying and resolving under-recognized shortcomings of existing procedures that frequently arise in scientific and industrial applications.
@@ 1 Introduction
Auto-regressive moving average (ARMA) models are the most well known and frequently used approach to modeling time series data. The general ARMA model was first described by Whittle [1], and popularized by Box and Jenkins [2]. Today, ARMA models are a fundamental component of various academic curricula, leading to their widespread use in both academia and industry. ARMA models are as foundational to time series analysis as linear models are to regression analysis, and they are often used in conjunction for regression with ARMA errors. A Google Scholar search for articles from 2024 onward that include the phrase “time series” and the term “ARMA” (or variants) yields over 18,000 results. While not all these articles focus on the same models discussed here, the importance of this model class to modern science cannot be overstated. Given the ubiquity of ARMA models, even small improvements in parameter estimation constitute a significant advancement of statistical practice.
@@
Imperfect likelihood optimization has an immediate consequence of complicating standard model selection procedures. Algorithms in widespread use lead to frequent inconsistencies in which a smaller model is found to have a higher maximized likelihood than a larger model within which it is nested. This is mathematically impossible but occurs in practice when the likelihood is imperfectly maximized, and is commonly observed using contemporary methods for ARMA models. Such inconsistencies are a distraction for the interpretation of results even when they do not substantially change the conclusion of the data analysis. Removing numeric inconsistencies can, and should, increase confidence in the correctness of statistical inferences.
There are various software implementations available for the estimation of ARMA model parameters. In this article, we focus on the standard implementations in R (stats package) and Python (statsmodels module), which we selected due to their widespread usage. While both implementations offer multiple ways to estimate parameters, the default approach in both software packages is to perform likelihood maximization, assuming that the error terms are Gaussian. The challenges in parameter estimation arise in this situation because there is no closed-form expression for the likelihood function, though computational algorithms do exist for maximizing ARMA likelihoods [3].
We begin by providing essential background information on the estimation of ARMA model parameters. We then present our proposed approach for parameter estimation, which leads to parameter values with a likelihood that is never lower and sometimes higher than the standard method. This is followed by a motivating example and a discussion of the potential implications of our proposed method. We also discuss the construction of standard errors for our maximum likelihood estimate. Specifically, we show that estimates of the standard error for model parameters that are default output of R and Python can be misleading, and we provide a reliable alternative. Throughout the article, we use the stats::arima function from the R programming language as a baseline for comparison. The same methodology for fitting parameters is used in the statsmodels.tsa module in Python, and we demonstrate that our results apply to this software as well (S2 Appendix).
2 Material and Methods
2.1 Maximum Likelihood for ARMA Models
Following the notation of Shumway and Stoffer [4], a time series $\{x_{t};t=0,\pm 1,\pm 2,\ldots\}$ is said to be $\text{ARMA}(p,q)$ if it is (weakly) stationary and
@@ 2.1 Maximum Likelihood for ARMA Models
with $\{w_{t};t=0,\pm 1,\pm 2,\ldots\}$ denoting a mean zero white noise (WN) processes with variance $\sigma_{w}^{2}>0$, and $\phi_{p}\neq 0$, $\theta_{q}\neq 0$. We refer to the positive integers $p$ and $q$ of Eq. 1 as the autoregressive (AR) and moving average (MA) orders, respectively. A non-zero intercept could also be added to Eq. 1, but for simplicity we assume that the time series $\{x_{t};t=0,\pm 1,\pm 2,\ldots\}$ has zero mean. We denote the set of all model parameters as $\psi=\{\phi_{1},\ldots,\phi_{p},\theta_{1},\ldots,\theta_{q},\sigma_{w}^{2}\}$. Our objective is to estimate model parameters using the observed uni-variate time series data.
Given the importance of ARMA models, numerous methods have been developed for parameter estimation. For instance, parameters can be estimated using Bayesian inference techniques [5, 6], neural networks [7], or by maximizing restricted likelihood [8, 9, 10], among other approaches. Specialized methods include those for integer-valued data [11], approaches robust to outliers [12], and non-Gaussian ARMA processes [13]. While each methodology has its merits, we focus on the approach most widely adopted in statistical practice: likelihood maximization, under the assumption that the WN process $\{w_{t}\}$ is Gaussian.
@@
A natural estimator would involve minimizing the sum of squares $\sum_{t=1}^{n}w_{t}^{2}$. However, since only $x_{1},x_{2},\ldots,x_{n}$ are observed and $w_{t}$ is recursively defined in Eq. 2 using values of $x_{t-p}$ and $w_{t-q}$, directly minimizing this sum is intractable. The CSS method addresses this issue by conditioning on the first $p$ values of the process, assuming $w_{p}=w_{p-1}=\ldots=w_{p+1-q}=0$, and minimizing the conditioned sum $\sum_{t=p+1}^{n}w_{t}^{2}$. While the CSS method provides an attractive solution due to its relative simplicity and easiness to compute, it ignores the error terms for the first few observations. This is particularly concerning when the time series is short or when there are missing observations. CSS minimization was previously popular because methods for likelihood maximization were considered prohibitively slow, though this is no longer the case with currently available hardware and software [14].
For likelihood maximization, Eq. 1 is reformulated as an equivalent state-space model. Although there are several ways this can be done, the approach of [3] is widely used. In this approach, we let $r=\max(p,q+1)$ and extend the set of parameters so that $\bar{\psi}=\{\phi_{1},\allowbreak\ldots,\allowbreak\phi_{r},\allowbreak\theta_{1},\allowbreak\ldots,\allowbreak\theta_{r-1},\allowbreak\sigma^{2}_{w}\}$, with some of the $\phi_{i}s$ or $\theta_{i}s$ being equal to zero unless $p=q+1$. We define a latent state vector $z_{t}\in\mathbb{R}^{r}$, along with transition matrices $T\in\mathbb{R}^{r\times r}$ and $Q\in\mathbb{R}^{r\times 1}$, enabling the recovery of the original sequence $\{x_{t}\}$ using Equations 3 and 4. For a detailed explanation of how to define the latent state $z_{t}$ and transition matrices $T$ and $Q$ in order to recover the ARMA model, we refer readers to Chapter 3 of Durbin and Koopman [15]. Along with initializations for the mean and variance of $z_{0}$, these equations allow for the exact computation of the likelihood of the ARMA model via the Kalman filter [16], which can subsequently be optimized by a numeric procedure such as the BFGS algorithm [17].
@@
Numeric black-box optimizers require an initial guess for parameter values. For ARMA models, this task is non-trivial because the valid parameter region is defined in terms of the roots of a polynomial associated with the parameters, as discussed in the next section. The default strategy in R and Python is to use the CSS estimator for initialization. This is an effective approach because the CSS estimator asymptotically converges to the MLE [4], and may therefore be close to the global maximum when there are sufficiently many observations. However, the CSS initialization is less useful with limited data, or when there are missing observations. The CSS estimate may also lie outside the valid parameter region, and in such cases, parameters are reinitialized at the origin. Both software implementations also allow for manual selection of initial parameter values, but finding suitable initializations manually can be challenging due to complex parameter inter-dependencies.
@@
The log-likelihood function of ARMA models is often multimodal [14], and therefore this single initialization approach can result in parameter estimates corresponding to local maxima (see Fig 1). This is true even for a carefully chosen initialization, such as the CSS estimate. A common strategy to optimize multimodal loss functions is to perform multiple optimizations using different initial parameters. However, we have found no instances of practitioners using a multiple initialization strategy for estimating ARMA model parameters. This may be explained by a general unawareness of the possibility of converging to a local maximum or because obtaining a suitable collection of initializations for ARMA models is nontrivial. For example, independently initializing parameters at random can place the parameter vector outside the region of interest. Furthermore, uniform random sampling generally fails to adequately cover the plausible parameter region (S1 Appendix).
@@
Fig 1: The profile log-likelihood of data simulated from four distinct MA(1) models, demonstrating a few examples of multimodal likelihood surfaces. The solid, black line indicates the true value of $\theta_{1}$; the dotted line is the CSS-initialization. The dashed lines correspond to the estimate $\hat{\theta}_{1}$ using stats:arima (red) and our proposed algorithm (implemented in arima2::arima, blue).
2.2 Novel Multi-start Algorithms
@@
To obtain random parameter initializations, parameter sets must correspond to causal and invertible ARMA processes; definitions are in Chapter 3 of Shumway and Stoffer [4]. Let $\{\phi_{i}\}_{i=1}^{p}$ and $\{\theta_{i}\}_{i=1}^{q}$ be the coefficients of the $\text{ARMA}(p,q)$ model (Eq. 1), and define $\Phi(x)=1-\phi_{1}x-\phi_{2}x^{2}-\ldots-\phi_{p}x^{p}$ as the AR polynomial, and $\Theta(x)=1+\theta_{1}x+\theta_{2}x^{2}+\ldots+\theta_{q}x^{q}$ as the MA polynomial. An ARMA model is causal and invertible if the roots of the AR and MA polynomials lie outside the complex unit circle. As a result, a direct approach to obtaining random parameter initilizations for the target parameter space is to sample the roots of $\Phi(x)$ and $\Theta(x)$ and use these values to reconstruct parameter initializations (Algorithm 1). Alternatively, vectors of length $p$ and $q$ that have elements in the interval (1,1)(-1,1) can be sampled, and then transformed into the parameter space using the Durbin-Levinson recursion (Algorithm 2), as described in [18].
7Get preliminary estimate: $\hat{\psi}_{0}=f(\psi_{0})$; set $k=0$;
@@
11     Sample paired roots as real with probability $p=\sqrt{1/2}$;